During a family gathering last christmas, Angel was able to collect 12 monetary bills consisting of 20 pesos bills and 50 pesos bills from her tito's and tita's. She received a total of 390 pesos from them. How many of each bill did Angel received?
Answer by ikleyn(52748)

.

Let x be the number of the 50-pesos banknotes.


Then the number of 20-pesos banknotes is (12-x).


The "money' equation is


    50x + 20*(12-x) = 390   pesos.


Simplify and solve for x


    50x + 240 - 20x = 390

    30x             = 390 - 240

    30x             = 150

      x             = 150/30 = 5.


ANSWER.  5 50-pesos banknotes and the rest,  (12-5) = 7 20-pesos banknotes.


CHECK.   5*50 + 7*20 = 250 + 140 = 390 pesos.    ! Precisely correct !

Solved.

-------------------

To see many similar solved problems, look into the lesson
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive
    - Three methods for solving standard (typical) coin word problems
in this site.

You will find there the lessons for all levels - from introductory to advanced,
and for all methods used - from one equation to two equations and even without equations.

A convenient place to quickly observe these lessons from a  "bird flight height"  (a top view)  is the last lesson in the list.

Read them and become an expert in solution of coin problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Coin problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.